• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 18:25:00 on February15,2008

    Entry number 228565

    keyword=Pi0 Work today

    Scott and Tony began work on moving the low voltage supplies and patch panels 
    to their own rack (moved from VME rack).
    They also removed the faulty HV power supplies from the DAQ weldment.  Those 
    were labeled and placed by the BigBite Linux machines.. so Jack knows which 
    ones they are.
    John and I moved a few of the baskets containing NaI electronics from the left 
    side to the right side of the beamline.
    Khem and I ran a selected bunch of cables (ribbon, low voltage, HV, signal) 
    from the DAQ weldment to the area underneath the BigBite stand.  All appear to 
    be of adequate length.
    Brad setup the Argon-CO2 gas for the hadron stack drift chambers.